# Trustworthy Pedestrian Trajectory Prediction via Pattern-Aware Interaction Modeling

### v2

@@ IV-B SSOS: Seq-Start of Seq
where $\widehat{pos}_{1:\tau}$ denotes the reconstructed observed trajectory from time $1$ to $\tau$; $\widehat{pos}_{\tau+1:T-1}$ denotes the predicted trajectory excluding the sampled goal; $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that balance the contributions of the reconstruction loss and the prediction loss; MSE is the Mean Squared Error.
However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction. These byproducts may introduce additional computational overhead.
IV-C Seq-CVAE: LSTM-Based CVAE
RNN (typically LSTM [20]) and Conditional Variational Autoencoder (CVAE) [36] have been widely applied in pedestrian trajectory prediction [37, 21, 6, 3]. We combine the two approaches and propose a generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals. The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$. These goals are subsequently fed into the Interaction Encoder to achieve its goal-driven functionality. Our purpose in this section is to learn a generative model $g_{\phi}(\widehat{pos}_{T}|c)$, where $\phi$} represents the model parameters. This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$. The Seq-CVAE architecture is illustrated in Figure 4.
@@ IV-C Seq-CVAE: LSTM-Based CVAE
$c_{Encoder}$ $=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\}))$ (4)
$c_{Decoder}$ $=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (5)
@@
V Experiment
TABLE I: Results comparison on all datasets. Bold: Best
@@ V Experiment
Datasets. We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [12] and UCY [13]. The datasets include five sub-datasets, covering four different scenarios and 1536 trajectories. These trajectories include various interaction scenarios such as walking together, as couples, and in groups. Similar to previous work [43, 6, 34], we adopt the leave-one-out strategy for splitting the training and testing sets. That is, four datasets are used for training and validation, and one dataset is used for testing. Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters. Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
Evaluation Protocol. Like many previous work [21, 8, 34], visual information is not used in our approach. During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds. For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choosing the best result for metric computation. This is the standard protocol in pedestrian trajectory prediction [40, 6, 7, 8].
Metrics. We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics. ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$, and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|$ (6)
$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory; $\widehat{pos}_{t}$ and $pos_{t}$ represent the predicted coordinates and ground truth coordinates at time step $t$, respectively.
Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section V-B.
Implementation Details. In our model, the Transformer component adopts the original Transformer architecture [35]. For the Neighbor Encoder module, the hidden size is set to 128. In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 4. In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance. Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2. Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation. To ensure the stability of the data distribution, all trajectories are shifted so that they start at coordinate (0,0). For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away. These values are subsequently normalized using Min-Max normalization.
During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently, and the teacher forcing technique is applied to facilitate training. For the Encoder-Generator, the SSOS strategy is applied. The learning rate is set to $1\times 10^{-4}$. For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$. The model is optimized using the Adam optimizer with 50 epochs. Mean Squared Error (MSE) serves as the loss function. During testing, the model generates predictions autoregressively. All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
V-A Comparison
V-A1 Comparison Baselines
We compare our model with representative methods PECNet [7] and TF [11], as well as the latest work [42, 4, 5, 9]. Additionally, we include some common baselines such as SoPhie [39] and CGNS [40]. The baseline results are from officially reported metrics.
V-A2 Evaluation
The comparative results are demonstrated in Table I. Compared to the aforementioned methods, our model achieves the best average ADE of 0.26. For goal prediction, the Seq-CVAE achieves excellent performance with an average FDE of 0.52, slightly trailing only behind PECNet [7] and TDAGCN [4] despite its relatively lightweight design. In contrast to FDE, which only evaluates the final predicted position, ADE measures error across all predicted trajectory points, with greater emphasis on intermediate trajectory deviations, thereby better reflecting the model’s capacity to model complex pedestrian interactions. To focus on the model’s ability to capture interactions, we adopt ADE as the evaluation metric for assessing and comparing performance across different scenarios.
@@ V-A2 Evaluation
Sparse Scenarios. However, on the Hotel dataset, our model underperforms relatively compared to the original Transformer [11]. The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning. Moreover, the dataset contains predominantly linear trajectories with sparse interactions. In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors. This issue is also shared by other complex models such as [4] and [42], as shown in Table I.
@@
V-B Ablation Studies
V-B1 Interaction-Related Components Analysis
For the interaction part, the region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects. The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor. To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table II. These variations allow us to isolate the individual and combined effects of region partition and interaction state. The original InSyn incorporates both region partition and interaction state. Case studies on the variations are provided in the supplementary material.
From the results in Table  II, it is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance. This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability. Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance. This would underperform when the training and testing sets differ. For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
TABLE II: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
@@ V-B1 Interaction-Related Components Analysis
V-B2 SSOS Strategy Analysis
Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper. SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks, such as pedestrian trajectory prediction. In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted. We replace the traditional single SOS with the sequence-based SSOS. This approach incorporates the observed values from steps $0:\tau$ as inputs for both the encoder and decoder during the initial stage. During the autoregressive process, the outputs are progressively concatenated and fed back into the Trajectory Generator as inputs. This allows the model to predict the complete trajectory while computing the loss for all outputs (Equation 3). To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
@@ V-B2 SSOS Strategy Analysis
Table III shows the average ADE results across all datasets, comparing different ablations of our model. ”w/o” refers to ”without” a component. The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components. The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy. This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories. Additionally, it mitigates error accumulation during autoregressive prediction.
TABLE III: IDE Result of Ablation Study on SSOS
@@
V-C Case Study
Figure 5: Case Study of Region Partition and Interaction State. This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model incorporating both components (InSyn). To isolate the effect of interaction modeling, we exclude the Goal Sampler in this evaluation.
To further investigate the effectiveness of region partition and interaction state, we conduct a Case study. The cases (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
In Sync Scenarios. Cases (a), (b), (e), and (f) represent instances where pedestrians walk in coordination with others nearby. In cases (a) and (b), where the target pedestrian moves alongside a companion, InSyn preserves the direction and velocity of the agent’s motion at the end of the observation period, closely matching the ground truth. In contrast, both ablated variants produce noticeable deviations. These cases suggest that the influence of the neighboring pedestrian is relatively weak. InSyn effectively models the behaviors by accurately capturing the reduced impact of the neighbor on the agent’s motion. Case (e) depicts a group walking situation with multiple neighboring pedestrians moving in sync. Here, only InSyn predicts the correct direction of motion, while both variants mistakenly infer an opposite trajectory, highlighting their failure to capture the low-impact nature of synchronized group behavior. Case (f) presents a simple, slow-moving scenario where all models achieve comparable performance, showing that the base model alone suffices for simple prediction tasks.
Conflict Scenarios. Cases (c) and (d) illustrate interactions where conflict plays a dominant role. In case (c), although the agent’s last observed motion trends slightly upward, InSyn predicts a flatter, rightward trajectory that better aligns with the ground truth. This adjustment reflects the model’s ability to recognize a conflicting pedestrian approaching from above, thereby reducing unnecessary upward movement. The w/o-RP and w/o-IS variants, lacking directional awareness and interaction-type recognition respectively, simply extrapolate the upward motion and thus produce larger errors. Case (d), however, reveals a limitation of InSyn. In this instance, the model appears to overestimate the influence of a pedestrian approaching from below, leading to an excessive upward deviation from the true path. This suggests that while InSyn improves reliability and interpretability in most complex scenarios, further refinement is needed to enhance control precision in handling conflict interactions.
@@ V-C Case Study
VI Conclusion
This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture pedestrian complex interaction patterns. By introducing interpretable structure into otherwise black-box models, InSyn improves interpretability and robustness in socially complex environments. Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more reliable predictions. Besides, our proposed SSOS strategy enhances prediction accuracy by mitigating initial-step divergence in numerical sequence modeling. A detailed case study further illustrates how modeling explicit interaction patterns contributes to interpretable and trustworthy predictions in diverse real-world scenarios.
Future work will focus on two key directions. First, we aim to design a more refined approach for partitioning interaction regions to better handle challenges posed by small-scale scenes. Second, we plan to apply the SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.
Acknowledgment
This work was supported by the Natural Science Foundation of Shanghai (No. 23ZR1422800). The authors would like to thank the SusCom Lab and Dr. Taowen Wang at HKUST(GZ) for their valuable support.

### v3

@@ 4.2 Trajectory Generator
where $\widehat{pos}_{1:\tau}$ denotes the reconstructed observed trajectory from time $1$ to $\tau$; $\widehat{pos}_{\tau+1:T-1}$ denotes the predicted trajectory excluding the sampled goal; $\lambda_{1}$ and $\lambda_{2}$ are used for balancing.
Ablation results in Section 5.2 shows that this strategy enchances the performance of the model. However, the SSOS strategy generates additional outputs $\widehat{pos}_{1:\tau}$, which are not part of the target prediction. These byproducts may introduce additional computational overhead.
4.3 Seq-CVAE Goal Sampler
RNN (typically LSTM [9]) and Conditional Variational Autoencoder (CVAE) [25, 12] have been widely applied in pedestrian trajectory prediction [37, 40, 1]. We combine the two approaches and propose a sequence-based generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals. The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$. These goals are subsequently fed into the Interaction Encoder to achieve its goal-driven functionality. Our purpose in this section is to learn a generative model $g_{\phi}(\widehat{pos}_{T}|c)$, where $\phi$} represents the model parameters. This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$. The Seq-CVAE architecture is illustrated in Figure 4.
@@ 4.3 Seq-CVAE Goal Sampler
$c_{Encoder}=\mathrm{\textit{FC}}(\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})),\quad c_{Decoder}=\mathrm{LSTM}(h_{\tau},\{pos_{0:\tau}\})$ (3)
@@
5 Experiment
5.0.1 Datasets and Evaluation Protocol.
We evaluate on the standard ETH [21] and UCY [14] datasets, following the leave-one-out strategy [30, 40, 39, 5]. As in prior work [37, 38, 39], we do not use visual information. The model observes 3.2 seconds (8 frames at 2.5 fps) and predicts the next 4.8 seconds. Evaluation employs the standard best-of-$K=20$ protocol [15, 40, 6, 38] for fair comparison.
5.0.2 Metrics.
We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics.
$ADE=\frac{1}{T-\tau}\sum_{t=\tau+1}^{T}\left\|\widehat{pos}_{t}-pos_{t}\right\|,\quad FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (4)
where $\tau$ represents the last time step of the observed trajectory; $T$ represents the last time step of the predicted trajectory;
Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE), which will be discussed in Section 5.2.
5.0.3 Implementation Details.
Our model builds on the Transformer architecture [28]. Key configurations: Neighbor Encoder hidden size 128; Seq-CVAE LSTM and latent sizes 256 (see Fig. 4); neighbor radius 2m. We apply data augmentation (rotation and scaling) and normalize trajectories to start at (0,0). The Interaction Encoder/Generator ($lr=1\times 10^{-4}$) and Seq-CVAE ($lr=1\times 10^{-3}$, KL weight=5) are trained separately for 500 epochs with Adam and MSE loss. Testing process is autoregressive. We performed comprehensive data processing and augmentation; the detailed pipeline is provided in our code.
5.1 Comparison
5.1.1 Comparison Baselines.
We compare our model with representative methods PECNet [6] and TF [7], as well as the rencent work [32, 31, 23, 17]. Additionally, we include some common baselines such as SoPhie [22] and CGNS [15]. The baseline results are from officially reported metrics.
5.1.2 Evaluation.
Table 1: Results comparison on all datasets. Bold: Best
@@ 5.1 Comparison
The comparative results are demonstrated in Table 1. Compared to the aforementioned methods, our model achieves the best average ADE of 0.26. For goal prediction, the Seq-CVAE achieves an average FDE of 0.52, which is competitive though slightly higher than some state-of-the-art methods [6, 31]. This is likely due to its separated design from the main interaction parts. Since ADE measures overall trajectory consistency and better reflects the model’s capacity to capture complex pedestrian interactions, we emphasize it as the primary metric for evaluation of interaction modeling.
@@
Sparse Scenarios. However, on the Hotel dataset, our model underperforms relatively compared to the original Transformer [7]. The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning. Moreover, the dataset contains predominantly linear trajectories with sparse interactions. In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors. This issue is also shared by other complex models such as [31] and [32], as shown in Table 1.
@@
5.2 Ablation Study
5.2.1 Interaction-Related Components Analysis.
For the interaction part, the region partition divides the area centered on the agent into four regions to capture directional interaction effects. The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor. To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table 2. “w/o” refers to “without” a component. The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components. These variations allow us to isolate the individual and combined effects of region partition and interaction state.
Table 2: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
@@ 5.2 Ablation Study
It is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance. This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability. Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance and direction. This would underperform when the training and testing sets differ. For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations. More case studies on the variations are provided in Section 5.3.
5.2.2 SSOS Strategy Analysis.
Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper. SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks (detail in Section 4.2), such as pedestrian trajectory prediction. To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
@@
The results in Table 3 show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy. This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories. Note that the improvement is more significant in Zara datasets, where pedestrians have fewer interactions and rely more on their own historical trajectories. The SSOS strategy, which reinforces the model’s encoding of individual motion patterns by integrating the observed sequence, thereby enhances self-expression and leads to greater gains.
Table 3: IDE Result of Ablation Study on SSOS
@@
5.3 Case Study
Figure 5: Case Study of Region Partition and Interaction State. This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model (InSyn). To isolate the effect of interaction modeling, we exclude the Goal Sampler in this evaluation.
To further investigate the effectiveness of region partition and interaction state, we conduct a case study. The cases extracted from the datasets (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
5.3.1 In Sync Scenarios.
Cases (a), (b), (e), and (f) represent instances where pedestrians walk in coordination with others nearby. In cases (a) and (b), where the target pedestrian moves alongside a companion, InSyn preserves the direction and velocity of the agent’s motion at the end of the observation period, closely matching the ground truth. In contrast, both ablated variants produce noticeable deviations. These cases suggest that the influence of the neighboring pedestrian is relatively weak. InSyn effectively models the behaviors by accurately capturing the reduced impact of the neighbor on the agent’s motion. Case (e) depicts a group walking situation with multiple neighboring pedestrians moving in sync. Here, only InSyn predicts the correct direction of motion, while both variants mistakenly infer an opposite trajectory, highlighting their failure to capture the low-impact nature of synchronized group behavior. Case (f) presents a simple, slow-moving scenario where all models achieve comparable performance, showing that the base model alone suffices for simple prediction tasks.
5.3.2 Conflict Scenarios.
Cases (c) and (d) illustrate interactions where conflict plays a dominant role. In case (c), although the agent’s last observed motion trends slightly upward, InSyn predicts a flatter, rightward trajectory that better aligns with the ground truth. This adjustment reflects the model’s ability to recognize a conflicting pedestrian approaching from above, thereby reducing unnecessary upward movement. The w/o-RP and w/o-IS variants, lacking directional awareness and interaction-type recognition respectively, simply extrapolate the upward motion and thus produce larger errors. Case (d), however, reveals a limitation of InSyn. In this instance, the model appears to overestimate the influence of a pedestrian approaching from below, leading to an excessive upward deviation from the true path. This suggests that while InSyn improves reliability and interpretability in most complex scenarios, further refinement is needed to enhance control precision in handling conflict interactions.
@@ 5.3 Case Study
6 Conclusion
This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture and leverage pedestrian complex interaction patterns. By introducing interpretable structure into otherwise black-box models, InSyn improves both accuracy and interpretability. Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more accurate predictions in complex interaction scenarios. Besides, our proposed SSOS strategy enhances prediction accuracy by mitigating initial-step divergence in numerical sequence modeling. A detailed case study further illustrates how modeling explicit interaction patterns contributes to interpretable and trustworthy predictions in diverse real-world scenarios. Future work will focus on two key directions. First, we aim to design a more refined approach for interaction identification. Second, we plan to apply the SSOS strategy to other types of numerical sequential data beyond trajectory prediction, exploring its potential in diverse domains.